OpenStudy (anonymous):

parallel plate capacitor which consists of plates of area 1cm2, separated by 2mm, is connected to a 10-V battery. What is the electric field between the plates?

OpenStudy (sidsiddhartha):

c=E0*A/D=(8.85*10^-12*1*10^-4)/(2*10^-3)=4.425*10^-13 now q=cv=4.425*10^-13*10=4.425*10^-12 C